# Probing Superheavy Dark Matter with Exoplanets

### v1

@@ I Introduction
Detecting Dark Matter (DM) via non-gravitational interactions remains a central focus of ongoing searches for new physics beyond the Standard Model (SM). Terrestrial experiments and astrophysical observations have imposed stringent constraints on the parameter space of the prevailing Weakly Interacting Massive Particle (WIMP) model. While extensive efforts have been dedicated to theoretical models and experimental searches for light DM [1, 2, 3], DM candidates heavier than the weak scale present another compelling direction within the DM parameter space for future searches [4]. The traditional upper bound on the mass of thermal DM is $\mathcal{O}(100)\,{\rm GeV}$ set by the perturbative unitarity in their production processes [5]. Recent studies have proposed alternative cosmological mechanisms for the production of DM with masses beyond the unitarity bound [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Experimental searches for these superheavy DM candidates, spanning from $100\,{\rm TeV}$ to the Planck scale $M_{\rm pl}$, remain challenging due to the suppressed DM flux in direct detection experiments, necessitating an extremely large exposure that is beyond the reach of current ground-based detectors. This has motivated the exploration of superheavy DM using astrophysical objects as targets through scattering and capture processes, driven by the same DM-SM interactions probed in terrestrial experiments. To date, extensive studies have explored the capture of DM particles by various astrophysical bodies, leading to a range of consequences, including heating, cooling, Black Hole (BH) formation in the core, neutrino emission, alterations in the equation of state of compact objects and their mass-radius relation, as well as gravitational wave emission. These astrophysical entities include neutron stars[18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40], white dwarfs[41, 42, 43, 44, 45, 46, 47, 48, 49], the Sun[50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64], the Earth[65, 66, 67], the Moon[68, 69], and other planetary bodies[70, 71, 72, 73, 74, 75, 76, 77, 78, 76, 79].
@@ II.1 Gas Planet Model
To simplify the analysis, we adopt an isentropic model, assuming constant entropy throughout the interiors of exoplanets, thereby avoiding complexities associated with entropy variations across different radii. We set the specific entropy per electron to $S=7\,k_{\mathrm{B}}\,\mathrm{e^{-1}}$ (or equivalently, $S\approx 7.61\,k_{\mathrm{B}}\,\mathrm{atom^{-1}}\approx 0.05\,\mathrm{MJ}\,\mathrm{kg^{-1}}\,\mathrm{K^{-1}}$), characteristic of Jupiter at its current age [103, 104, 91]. Here $k_{\mathrm{B}}$ is the Boltzmann constant. More complex models of Jupiter, featuring varying entropy at different radii, have been studied in [91, 94, 93]. However, deviations from constant entropy do not significantly affect the final results of this work. We show the exoplanet model in Fig. 1. The left panel shows the temperature-density relationship of the hydrogen-helium mixture corresponding to constant entropy values. For comparison, we present different benchmark values: $6\,k_{\mathrm{B}}\,\mathrm{e^{-1}}$ (green), $7\,k_{\mathrm{B}}\,\mathrm{e^{-1}}$ (blue), $8\,k_{\mathrm{B}}\,\mathrm{e^{-1}}$ (red), $10\,k_{\mathrm{B}}\,\mathrm{e^{-1}}$ (magenta), and $16\,k_{\mathrm{B}}\,\mathrm{e^{-1}}$ (brown), adopted from the aforementioned data. Also shown are the temperature-density contour for two objects with masses equal to $1\,M_{\rm J}$ (solid black) and $10~{}M_{\rm J}$ (dashed black), as presented in [103, 104], for comparison with the isentropic model. As illustrated, the Jupiter-mass object closely aligns with the isentropic contour of $7\,k_{\mathrm{B}}\,\mathrm{e^{-1}}$, while the $10M_{\mathrm{J}}$ object aligns with the isentropic contour of $8\,k_{\mathrm{B}}\,\mathrm{e^{-1}}$.
@@ III Dark Matter Capture in Exoplanets
The $\mu_{j}$ and $\mu_{\mathrm{p}}$ are the reduced masses of element ${j}$ and nucleon with $m_{\mathrm{\chi}}$, respectively, and $\sigma_{\mathrm{\chi p}}$ is the scattering cross section of the DM particle with the nucleon. The second model assumes that the scattering cross section from the nucleus is independent of the mass number as
@@ IV Black Hole Formation in Exoplanets
We note that this represents a conservative assumption for the BH formation process, as the time scales in Eq. (11) may overlap when multiple processes occur simultaneously. The $t_{\rm cap}$ is the time required to capture a sufficient DM abundance. Given the capture rate from the last section, the capture time is expressed as $t_{\rm cap}=M_{\rm crit}/(m_{\mathrm{\chi}}C)$, where $M_{\rm crit}$ is the critical mass required for BH formation, which will be introduced later in this section. The other time scales, $t_{\rm th}^{\rm I}$, $t_{\rm th}^{\rm II}$, $t_{\rm drift}$, and $t_{\rm col}$, correspond to the first and second thermalization times, the drift time, and the collapse time, respectively, which will be discussed in detail below.
@@ IV.3 Collapse Time
[…]
@@ V Results
The detection of planetary-mass BHs greatly benefits from existing exoplanet detection methods. In the case of exoplanets with host stars, their transmutation into BHs can still influence the motion of the host stars and be detected using Doppler spectroscopy and astrometric methods. However, since the geometric size of a BH is too small to produce observable effects in the star’s light curve, the number of transit events will decrease if exoplanets transition into BHs due to DM capture. The impact of superheavy DM can be assessed by comparing the number of observed transit events with the number of detections through Doppler spectroscopy or astrometric methods, similar to the approach proposed in [141, 142] for detecting primordial BHs and dark exoplanets in extrasolar systems. While a single event cannot distinguish DM-induced BH formation from other scenarios, the large future exoplanet database may help identify it through population-level mass distributions. A notable number of missing transit events is expected in systems identified through stellar motion observations, particularly for higher-mass planets, where BH formation is more efficient. Furthermore, this reduction is anticipated to be more significant in systems closer to GC, where the DM energy density is higher.
@@
The microlensing signals, however, appear similar for both exoplanets and BHs of the same mass due to the universality of gravitation in the point-lens limit. Since microlensing is less sensitive to the density profiles of detected objects, follow-up analysis of light curve features is essential to distinguish BHs from exoplanets of similar mass. Future improvements can be incorporated into the lensing event analyses in this direction. In [143], the distribution of lens candidates is analyzed in the two-dimensional parameter space of the Einstein timescale and microlensing parallax, aiming to distinguish BHs from other compact objects at stellar masses. However, differentiating between other classes of lens remains challenging with current techniques. The upcoming Roman telescope will be capable of measuring the astrometric information of lenses, providing the possibility of a three-dimensional parameter space to distinguish transmuted BHs from original rogue exoplanets. In addition to identification at the single-event level, the mass distribution of planetary lenses, such as the galactic population of rogue planets recently studied in [144], will help establish the expected event rate at a population level, providing a basis for comparison with future observations.
@@ Appendix A Dark Matter Velocity Distribution
[…]

### v2

@@ I Introduction
Detecting Dark Matter (DM) via non-gravitational interactions remains a central focus of ongoing searches for new physics beyond the Standard Model (SM). Terrestrial experiments and astrophysical observations have imposed stringent constraints on the parameter space of the prevailing Weakly Interacting Massive Particle (WIMP) model. While extensive efforts have been dedicated to theoretical models and experimental searches for light DM [1, 2, 3], DM candidates heavier than the weak scale present another compelling direction within the DM parameter space for future searches [4]. The traditional upper bound on the mass of thermal DM is $\mathcal{O}(100)\,{\rm GeV}$ set by the perturbative unitarity in their production processes [5]. Recent studies have proposed alternative cosmological mechanisms for the production of DM with masses beyond the unitarity bound [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Experimental searches for these superheavy DM candidates, spanning from $100\,{\rm TeV}$ to the Planck scale $M_{\rm pl}$, remain challenging due to the suppressed DM flux in direct detection experiments, necessitating an extremely large exposure that is beyond the reach of current ground-based detectors. This has motivated the exploration of superheavy DM using astrophysical objects as targets through scattering and capture processes, driven by the same DM-SM interactions probed in terrestrial experiments. To date, extensive studies have explored the capture of DM particles by various astrophysical bodies, leading to a range of consequences, including heating, cooling, Black Hole (BH) formation in the core, neutrino emission, alterations in the equation of state of compact objects and their mass-radius relation, as well as gravitational wave emission. These astrophysical entities include neutron stars[18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45], white dwarfs[46, 47, 48, 49, 50, 51, 52, 53, 54, 55], the Sun[56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71], the Earth[72, 73, 74], the Moon[75, 76], and other planetary bodies[77, 78, 79, 80, 81, 82, 83, 84, 85, 83, 86].
@@ II.1 Gas Planet Model
To simplify the analysis, we adopt an isentropic model, assuming constant entropy throughout the interiors of exoplanets, thereby avoiding complexities associated with entropy variations across different radii. We set the specific entropy per electron to $S=7\,k_{\mathrm{B}}\,\mathrm{e^{-1}}$ (or equivalently, $S\simeq 7.61\,k_{\mathrm{B}}\,\mathrm{atom^{-1}}\simeq 0.05\,\mathrm{MJ}\,\mathrm{kg^{-1}}\,\mathrm{K^{-1}}$), characteristic of Jupiter at its current age [110, 111, 98]. Here $k_{\mathrm{B}}$ is the Boltzmann constant. More complex models of Jupiter, featuring varying entropy at different radii, have been studied in [98, 101, 100]. However, deviations from constant entropy do not significantly affect the final results of this work. We show the exoplanet model in Fig. 1. The left panel shows the temperature-density relationship of the hydrogen-helium mixture corresponding to constant entropy values. For comparison, we present different benchmark values: $6\,k_{\mathrm{B}}\,\mathrm{e^{-1}}$ (green), $7\,k_{\mathrm{B}}\,\mathrm{e^{-1}}$ (blue), $8\,k_{\mathrm{B}}\,\mathrm{e^{-1}}$ (red), $10\,k_{\mathrm{B}}\,\mathrm{e^{-1}}$ (magenta), and $16\,k_{\mathrm{B}}\,\mathrm{e^{-1}}$ (brown), adopted from the aforementioned data. Also shown are the temperature-density contour for two objects with masses equal to $1\,M_{\rm J}$ (solid black) and $10~M_{\rm J}$ (dashed black), as presented in [110, 111], for comparison with the isentropic model. As illustrated, the Jupiter-mass object closely aligns with the isentropic contour of $7\,k_{\mathrm{B}}\,\mathrm{e^{-1}}$, while the $10M_{\mathrm{J}}$ object aligns with the isentropic contour of $8\,k_{\mathrm{B}}\,\mathrm{e^{-1}}$.
@@ III Dark Matter Capture in Exoplanets
The $\mu_{j}$ and $\mu_{\mathrm{p}}$ are the reduced masses of element j{j} and nucleon with $m_{\mathrm{\chi}}$, respectively, and $\sigma_{\mathrm{\chi p}}$ is the scattering cross section of the DM particle with the nucleon. The second model assumes that the scattering cross section from the nucleus is independent of the mass number as
@@ IV Black Hole Formation in Exoplanets
We note that this represents a conservative assumption for the BH formation process, as the time scales in Eq. (11) may overlap when multiple processes occur simultaneously. The $t_{\rm cap}$ is the time required to capture a sufficient DM abundance. Given the capture rate from the last section, the capture time is expressed as $t_{\rm cap}=M_{\rm crit}/(m_{\mathrm{\chi}}C)$, where $M_{\rm crit}$ is the critical mass required for BH formation, which will be introduced later in this section. The other time scales, $_{\rm th}^{\rm I}, tt$$_{\rm th}^{\rm II}$, $t_{\rm drift}$, and $t_{\rm col}$, correspond to the first and second thermalization times, the drift time, and the collapse time, respectively, which, following [74], will be discussed in detail below.
@@ IV.3 Collapse Time
We calculate the time scale gravitational collapse in the following. In contrast to the second thermalization process, a DM particle initially undergoes gravitational infall in the viscous regime as its initial virialized velocity, determined with $K_{\rm i}=GM_{\rm crit}m_{\mathrm{\chi}}/2r_{\rm th}$, is smaller than that of the baryonic particles. As the DM particle is accelerated during the infall, it eventually gains sufficient kinetic energy to transition into the inertial regime. Since the DM velocity is much larger in the inertial regime, the collapse during this stage completes rapidly. Therefore, the end of the collapse process is effectively determined by the transition from the viscous to the inertial regime, indicated by $v_{\mathrm{\chi}}=v=\sqrt{3T\left(0\right)/\bar{m}}$, or equivalently, when the kinetic energy reaches $K_{\rm f}=3T\left(0\right)m_{\chi}/2\bar{m}$. The changes in kinetic energy, $\Delta K=K_{\rm f}-K_{\rm i}$, of the virialized system equals the total energy loss caused by viscous drag. Therefore, one can use the energy loss rate in the viscous regime in Eq. (16) to determine the collapse time with $\int_{0}^{t_{\rm col}}\,dE=\Delta K$. The collapse time is given by
@@ V Results
The detection of planetary-mass BHs greatly benefits from existing exoplanet detection methods. In the case of exoplanets with host stars, their transmutation into BHs can still influence the motion of the host stars and be detected using Doppler spectroscopy and astrometric methods. However, since the geometric size of a BH is too small to produce observable effects in the star’s light curve, the number of transit events will decrease if exoplanets transition into BHs due to DM capture. The impact of superheavy DM can be assessed by comparing the number of observed transit events with the number of detections through Doppler spectroscopy or astrometric methods, similar to the approach proposed in [148, 149] for detecting primordial black holes (PBHs) and dark exoplanets in extrasolar systems. While a single event cannot distinguish DM-induced BH formation from other scenarios, the large future exoplanet database may help identify it through population-level mass distributions. A notable number of missing transit events is expected in systems identified through stellar motion observations, particularly for higher-mass planets, where BH formation is more efficient. Furthermore, this reduction is anticipated to be more significant in systems closer to GC, where the DM energy density is higher.
@@
The microlensing signals, however, appear similar for both exoplanets and BHs of the same mass due to the universality of gravitation in the point-lens limit. Since microlensing is less sensitive to the density profiles of detected objects, follow-up analysis of light curve features is essential to distinguish BHs from exoplanets of similar mass. Future improvements can be incorporated into the lensing event analyses in this direction. In [150], the distribution of lens candidates is analyzed in the two-dimensional parameter space of the Einstein timescale and microlensing parallax, aiming to distinguish BHs from other compact objects at stellar masses. However, differentiating between other classes of lenses remains challenging with current techniques. The upcoming Roman telescope will be capable of measuring the astrometric information of lenses, providing the possibility of a three-dimensional parameter space to distinguish transmuted BHs from original rogue exoplanets. In addition to identification at the single-event level, the mass distribution of planetary lenses, such as the galactic population of rogue planets recently studied in [151], will help establish the expected event rate at a population level, providing a basis for comparison with future observations.
Gravitational microlensing is a powerful tool for detecting planet-sized objects in the Milky Way. While long-period and free-floating Jupiter-mass exoplanets have been observed through infrared surveys [152, 153, 154], only microlensing techniques currently allow for the detection of Earth- to Mars-mass planets that are on wide orbits or are free-floating. This is because microlensing does not rely on the luminosity of the object of interest, but instead depends on the architecture of the planetary system, the masses of the objects in the system, and the geometry relative to a background bright lens object.
Microlensing surveys toward the Galactic bulge include the Optical Gravitational Lensing Experiment (OGLE) [155, 123], Microlensing Observations in Astrophysics (MOA) [156, 157], the WISE Microlensing Survey [158], and the Korea Microlensing Telescope Network (KMTNet) [159]. Beyond the Galactic bulge, surveys have also targeted the Magellanic Clouds, including MACHO/EROS [160, 161, 162, 163, 164] and the OGLE-III and OGLE-IV phases [165, 166]. In addition, microlensing observations of the Andromeda galaxy have been conducted using the Subaru Hyper Suprime-Cam (HSC) [167, 168, 169].
Microlensing surveys have been utilized to study the abundance of PBHs and to constrain the fraction ($f$) of dark matter composed of PBHs. In a recent study, Ref. [166] summarized existing microlensing bounds and extended them using data from the OGLE High-Cadence Survey of the Magellanic Clouds. Additionally, Ref. [170] compiled more comprehensive constraints beyond microlensing, including those from the gamma-ray background, supernova lensing, dynamical heating of dwarf galaxies, wide binary stars, X-ray binaries, CMB distortions from accreting plasma in early universe, and disk stability arguments. This reference also presented projected future constraints from upcoming experiments such as the Roman Space Telescope, the Rubin Observatory Legacy Survey of Space and Time (LSST), and a future MeV gamma-ray facility. Among existing bounds, the strongest constraint on $f$ for Jupiter-mass black holes is $f<10^{-3}$, while future projections suggest sensitivity down to $f<10^{-4}$.
In our work, Jupiter-sized exoplanets within the viable parameter space can transmute into black holes with masses approximately equal to that of Jupiter. To estimate the total mass of such transmuted black holes, we assume there is, at most, roughly one Jupiter-sized planet per star in the Milky Way (see Ref. [171] for estimates of the free-floating exoplanet population). Given that the Milky Way contains on the order of $10^{11}$ stars, the total mass of Jupiter-sized objects would be approximately $10^{8}\,M_{\odot}$. If these exoplanets were to fully transmute into black holes, and assuming the dark matter halo mass is on the order of $10^{12}\,M_{\odot}$, the resulting maximum mass fraction from this mechanism alone would be $f\sim 10^{-4}$, which is comparable to current and projected upper limits from microlensing constraints. Therefore, our scenario of Jupiter-mass black hole formation from exoplanet collapse is not excluded by existing bounds on PBHs. Future surveys will be particularly valuable, as they will probe with greater sensitivity and place stronger constraints on $f$. This will help further test the viability of the exoplanet-collapse scenario as a possible origin of Jupiter-mass black holes.
@@ Appendix A Dark Matter Velocity Distribution
$\sigma_{\rm v}(r)^{2}$ $=$ $\frac{1}{\rho_{\mathrm{\chi}}\left(r\right)}\int_{r}^{\infty}\rho_{\mathrm{\chi}}\left(r^{\prime}\right)\frac{dV_{\rm MW}\left(r^{\prime},0\right)}{dr^{\prime}}dr^{\prime}.$ (37)